# Bifurcations and multistability in inducible three-gene toggle switch networks

### v1

@@ Abstract
Control of transcription presides over a vast array of biological processes including through gene regulatory circuits that exhibit multistability. Two- and three-gene network motifs are often found to be critical parts of the repertoire of metabolic and developmental pathways. Theoretical models of these circuits, however, typically vary parameters such as dissociation constants, transcription rates, and degradation rates without specifying precisely how these parameters are controlled biologically. In this paper, we examine the role of effector molecules, which can alter the concentrations of the active transcription factors that control regulation, and are ubiquitous to regulatory processes across biological settings. We specifically consider allosteric regulation in the context of extending the standard bistable switch to three-gene networks, and explore the rich multistable dynamics exhibited in these architectures as a function of effector concentrations. We then study how the conditions required for tristability and more complex dynamics, and the bifurcations in dynamic phase space upon tuning effector concentrations, evolve under various interpretations of regulatory circuit mechanics, the underlying activity of inducers, and perturbations thereof. Notably, the biological mechanism by which we model effector control over dual-function proteins transforms not only the phenotypic trend of dynamic tuning but also the set of available dynamic regimes. In this way, we determine key parameters and regulatory features that drive phenotypic decisions, and offer an experimentally tunable structure for encoding inducible multistable behavior arising from both single and dual-function allosteric transcription factors.
@@ I Introduction
Biological processes rely on intricate networks of gene interactions that together encode a multitude of possible cellular functions. Often, these genes produce proteins, called transcription factors, that alter the expression of other genes. Whether considering resource consumption in E. coli [1] or stages in animal development such as digit formation [2, 3, 4], vulval development [5, 6], or stem cell differentiation [7, 8, 9], such biological systems can precisely tune the relevant gene interactions toward many possible distinct phenotypes. This inherent multistability is crucial to cell function, allowing signaling from the environment, from metabolic resource constraints, or from other sources to induce systems toward what at times can be dramatically different end states.
@@
Determining from experimental data alone the exact regulatory interactions that collectively drive multistable dynamics, however, is another matter. Indeed, the sensitivities of dynamics and steady state outcomes to both internal and external tuning are challenging to disentangle. To better understand these dynamics, it is useful to explore simpler motifs that establish the building blocks for multistability in higher-order networks.
@@
One regulatory motif that has emerged as an integral component of natural and synthetic genetic circuit design is the bistable toggle switch, in which each of the two genes produces a protein that represses expression of the other. Study of the lysis vs. lysogeny decision in bacteriophage lambda in the 1960-80s established the significance of this motif experimentally [10, 11, 12, 13], and subsequent research has modeled such motifs extensively using Hill models [14, 15, 16]. As a two-state model, however, the bistable switch sheds only partial light on the multistability of realistic phenotypic expression in cells.
Insights into the emergence of multistable dynamics in larger networks rely not only on the choice of model system but also on the approach to modeling these motifs. Existing literature on multistable cell fates during development, for example, has made significant strides by considering gene networks in terms of the Waddington landscape [17], and discusses the dynamic thresholds, or bifurcations, that can emerge from gene circuits described as attractor systems [18, 19, 20, 21]. When linking this interpretation to mathematical models, however, such studies often represent systems using either a set of network-free flow equations or a potential gradient. These approaches are undoubtedly useful for exploring dynamic stability through the tuning of theoretically accessible internal parameters such as dissociation constants, production rates, and degradation rates. In this paper, we directly link these properties to experimentally accessible tuning knobs, and precisely specify how the relevant parameters are quantitatively controlled in living cells [22].
@@
In fact, the activity of transcription factors often depends on the presence or absence of effector molecules [23, 24, 25, 26, 27, 28, 29, 30, 31, 22]. When targeting a repressor, for example, an effector could bind and stabilize a protein in its inactive configuration, which suppresses its function and induces expression. Certain transcription factors are also known to have context-dependent behavior, acting as repressors at one promoter site along the DNA and activators at another [32, 33]. Exploring different interpretations of the mechanistic role of effectors may then reveal critical differences in tuned expression that allow for complementary experimental investigation.
@@
We thus focus on examining the role of the effector concentration as a dynamical system parameter. Tuning inducer concentration alone, we can then directly observe corresponding shifts in dynamics. This framework offers a perspective for parametrization that is accessible to the experimentalist and also inherently to cell activity itself. For example, researchers can directly input external non-metabolizable effectors to study the dynamic shifts observed in theory [34], and cells can also naturally use one or more effectors, with two inducers driving cell fate determination in the lac operon and in cell differentiation during development, among other settings [35, 16]. Given the extensive experimental evidence emphasizing the significance of effectors as natural tuning knobs, we aim to theoretically examine how dynamics respond to explicit definitions for effectors and for the biological mechanism of induction.
@@
To analyze the roles of effector molecules in the emergence of multistability, we specifically study extensions of the toggle switch to three genes as shown in Figure 1. These motifs remain simple enough to visualize and physically interpret, yet also provide an additional layer of tunable gene interaction parameters that can allow more complex multistable dynamics – in this case, tristability and beyond. In exploring how a set of more than two genes couple to collectively decide among multiple possible cell fates, we gain insights into how these interactions drive dynamics in larger and more complex networks.
@@
The three-gene toggle switch provides a powerful basis for examining the emergence of tristability and higher-order multistability in a range of biological contexts. The motif is most directly useful in a synthetic setting, where dynamic shifts can be tuned precisely through experiments. However, the model can also help uncover how various sets of three (or more) key genes interact to generate multistable cell fates in nature. For example, experimental work has observed sets of three “master regulator” genes that interact to drive protein expression toward three or more distinct cell fate phenotypes, such as in T helper cell differentiation [36] and stem cell differentiation [37].
@@
Recent work has already begun to examine the comprehensive picture of stable expression in such three-gene toggles [38, 39]. Our study considers various approaches to induction in an alternative Hill model for the three-gene toggle. Within this reframed context for tuning dynamics, we then investigate not only how the mechanisms of induction shift transcription factor activities and stable state expressions, but also how the dynamic topologies of the resulting fixed point landscapes evolve.
@@
Figure 2: Expression of repressor 1 in the three-gene toggle switch. Equivalent definitions apply to expression of repressors 2 and 3. (A) Thermodynamic states, weights, and rates for expression of repressor 1. The regulating repressors ($R_{2}$ and $R_{3}$) bind non-exclusively to the target promoter region to suppress gene transcription. Each repressor $R_{i}$ binds at the promoter region for repressor $R_{1}$ with affinity $K_{1i}$. (B) Thermodynamic states, weights, and rates for expression of repressor 1 in the presence of inducers. Expression now depends on the active concentration of each repressor $R_{i}$, which is determined by a distinct inducer at concentration $c_{i}$, defining the probability of activity $p_{\text{act}}(c_{i})$.
We begin in Section II by defining our model for the three-gene toggle switch, and how to mathematically incorporate the regulatory activity of transcription factors as allosterically-inducible proteins. From this initial framework, Section III then demonstrates the different types of dynamics possible as inducer concentrations rise. In Section IV we highlight how the thresholds, i.e., bifurcations, that separate different dynamic regimes shift when perturbed away from an initial assumed symmetry in gene interactions. Section V explores the impacts of additional induction and self-activation on the complexity of dynamics and the number of stable steady states possible for the three-gene toggle switch. In the latter case in particular, we find that the biological interpretation of the mechanism for dual repression and self-activation, whether involving binding exclusivity or the effector’s role, can significantly affect the types of dynamics that unfold. In so doing, our work sheds light on the complexity of dynamics possible from simple changes in modeling choices, and on just how precisely inducer concentrations must be tuned under different conditions for a system to access various dynamic regimes. The diverse range of input-output responses demonstrated across the different three-gene networks presented here only further underscores the importance of understanding how these results generalize to higher-order gene regulatory architectures.
@@ II The three-gene toggle switch
We first consider the most direct extension of the two-gene bistable toggle switch to three genes, as depicted in Fig. 1. Transcribing a given gene $g_{i}$ generates a protein $R_{i}$ called a transcription factor, which can in turn bind to the promoter region of either remaining gene. In this particular incarnation of the model, the transcription factor prevents RNA polymerase (RNAP) from binding and initiating transcription, thereby repressing expression. We begin by discussing how to model the dynamics of this system of repressors before demonstrating the role of induction in governing how dynamical behavior evolves as a function of inducer concentration(s).
@@ II.1 The baseline model
This general model form assumes that the dynamics of gene transcription, i.e., the production of mRNA transcripts, evolves at a timescale such that production can be measured equivalently from the concentration of output repressor $R_{i}$. The first term of Eqn. 1 characterizes protein production at a maximal rate $a_{i}$, and depends on the probability $p_{i}^{\text{expr}}$ that expression can occur. The second term accounts for protein degradation at a rate $1/\tau_{i}$.
Fig. 2(A) breaks down the possible regulatory states and corresponding statistical weights that contribute to the probability of expression $p_{i}^{\text{expr}}$. This approach to describing biological regulation draws from well-established statistical mechanical principles [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. From this table, and equivalent definitions for expression of repressors 2 and 3, we define the dynamics of this three-gene circuit by the set of differential equations
@@
Each repressor $R_{j}$ binds non-exclusively with affinity $K_{ij}$ to the promoter region for the gene that expresses repressor $R_{i}$. This repression coefficient $K_{ij}$ thus reflects the characteristic concentration of repressor $R_{j}$ required to strongly regulate $R_{i}$’s expression. In other words, when $R_{j}=K_{ij}$, the produced concentration of repressor $R_{i}$ is half its maximum possible value.
Note that these equations describe transcription factor binding phenomenologically through a Hill function model with coefficient $n$ [55, 56]. This approach coarse-grains out the molecular detail of precise binding site occupancies found in thermodynamic models, instead representing the high cooperativity-limit of the thermodynamic model. The Hill coefficient $n$ thus does not necessarily correspond to the number of bound repressor molecules, but rather measures the sensitivity of output response to binding.
While there are advantages to the specificity of thermodynamic models [57], using them to evaluate stability profiles quickly becomes intractable in higher-dimensional networks, including those analyzed here. We therefore take a Hill model approach for the remainder of the discussion. Even without full knowledge of precise binding site occupancies, the Hill function offers a strong fit to empirical data for gene regulatory dynamics [14, 58]. The model also reflects the Hopfield barrier for the sharpness of input-output response [59], meaning that for a given coefficient $n$, the Hill model reflects the strongest, most sigmoidal input-output response possible for a system without energy expenditure. The Hill approach thus allows us to investigate dynamics under the most input-sensitive model conditions, and to observe its maximum capability for dynamic complexity.
We now have a model for the toggle switch that depends on various theoretically accessible parameters, specifically dissociation constants $K_{ij}$, production rates $a_{i}$, and degradation rates $1/\tau_{i}$. As pointed out previously, however, these parameters as written are not easily accessible experimentally, and are disconnected from the underlying biology that controls them within living cells. While the current model indicates that regulation is controlled by the total concentrations of transcription factors, these proteins are often themselves controlled by the allosteric binding of effector molecules. Cells therefore respond to changes in the number of active transcription factors.
@@ II.2 A statistical mechanical model for allosteric induction
Shifting to an allosteric description of gene regulation, we now argue that gene expression depends not on the total concentration of the regulating transcription factor, $\text{TF}_{\text{tot}}]$, but rather on the active concentration thereof, i.e.,
@@
where $p_{\text{act}}(c)$ is the probability that the transcription factor is active as a function of effector concentration $c$. This probability can be defined by the Monod-Wyman-Changeux (MWC) model [23, 24, 25, 26, 29, 30, 31, 22], which states that when an effector at concentration $c$ can bind allosterically at $m$ sites on a transcription factor, the probability that the transcription factor is active is
@@
Fig. 3(A) plots one possible probability curve $p_{\text{act}}(c)$ in the induction setting with $K_{I}<K_{A}$ for a fixed set of parameters. This function follows standard sigmoidal behavior, with a sufficient increase in inducer concentration rendering repressor inactive. Depending on how we define parameters, we can alter several key properties of the input-output response from this probability function. The leakiness represents the amount of activity at saturating concentrations of inducer, i.e., $p_{\text{act}}(c\rightarrow\infty)$, and the saturation conversely denotes the amount of activity in the absence of inducer ($p_{\text{act}}(c=0)$). The inducer concentration at which the probability is half maximal is denoted as the $EC_{50}$ value, and the sharpness of the input-output response curve at this inflection point increases with an increasing Hill coefficient $n$. The MWC parameters chosen therefore determine how effectively inducers suppress the activity of repressors and the sensitivity of the output response to an increase in inducer concentration.
We determine the biologically permissible ranges for these MWC parameters ($K_{I}$, $K_{A}$, and $\Delta\epsilon$) by making several assumptions. First, the absence of inducer does not imply the complete absence of inactive repressor, for there may be some natural fractional error $q$. We suppose, then, that in the absence of inducer,
$p_{\text{act}}(c=0)\geq 1-q,$ (7)
or alternatively,
Rearranging Eqn. 8, we obtain a condition for the energy difference $\Delta\epsilon$ between the inactive and active repressor configurations that
$e^{-\beta\Delta\epsilon}\leq\frac{q}{1-q}.$ (9)
It follows that for the probability of activity to perfectly saturate at unity in the absence of inducer, with $q=0$, the energy of the inactive repressor state must significantly exceed that of the active state so that $\Delta\epsilon\rightarrow\infty$ and the repressor is always active.
Gene expression is also generally “leaky,” meaning there is typically a basal level of expression even in the presence of a large concentration of repressor. We assume that this leakiness is also reflected by the fraction $q$ such that a fraction of repressors less than $q$ are active in the limit of infinite inducer concentration, with
$=\frac{(K_{I})^{m}}{(K_{I})^{m}+e^{-\beta\Delta\epsilon}(K_{A})^{m}}\leq q.$ (10)
Rearranging Eqn. 10, we obtain a second condition
Eqns. 9 and 11 thus establish biologically reasonable bounds on the dissociation constants $K_{A,\>I}$ and the energy difference $\Delta\epsilon$ for a given “error” $q$ and number of binding sites $m$.
@@
Figure 3: Activity of repressors as a function of inducer concentration $c$, for which $K_{I}<K_{A}$. (A) The probability of active repressor as a function of inducer concentration, with $m=2$, $\epsilon=4\>k_{B}T$, $K_{A}=150$ $\mu M$, and $K_{I}=5$ $\mu M$. The saturation and leakiness limits are denoted in blue and orange, respectively, and the $EC_{50}$, i.e., inducer concentration at which the probability is half maximal, is marked in purple. (B) Evolution of probability for different values of $K_{I}$, with the curve in panel (A) shown again in green. The shift in probability spans the range of allowable $K_{I}$ values for the specified parameter set up to the boundary at $K_{I}=6.77\>\mu M$, and assumes a fractional error of at most $q=0.1$ in saturation and leakiness.
Suppose, for example, that less than $10\%$ ($q=0.1$) of repressors are active in the presence of infinite inducer concentration. The parameters of this MWC model are then bounded by
$e^{-\beta\Delta\epsilon}\leq\frac{1}{9}$ (12)
and
$\Big(\frac{K_{A}}{K_{I}}\Big)^{m}\geq 9e^{\beta\Delta\epsilon}.$ (13)
@@
Fig. 3(B) highlights several viable probability curves within the bounded parameter regime defined by Eqns. 12 and 13. We assume throughout that there are $m=2$ binding sites for inducers on repressors. For fixed $\beta\Delta\epsilon=4$ and $K_{A}=150$ $\mu M$, Fig. 3(B) then maps the probability of activity as a function of molar inducer concentration for increasing dissociation constant $K_{I}<K_{A}$. As the ratio $K_{d}^{A}/K_{d}^{I}$ decreases, approaching its lower bound (denoted in cyan in Fig. 3), the curve’s inflection point shifts toward higher inducer concentrations. This occurs because as $K_{d}^{I}$ increases and the inducer becomes less tightly bound to the inactive compared to the active repressor, a higher inducer concentration is necessary to sequester inactive repressors equivalently. Given that the choice of parameters in Fig. 3(A) is also within the biologically reasonable regime, we use these same parameters for the remainder of this work, unless otherwise stated.
@@
Note that increasing the number of inducer binding sites $m$ for the target repressor protein would increase the cooperativity of inducer-repressor interactions, and thus increase the sensitivity of response to the presence of inducer. This would make the slope of the probability function steeper at the inflection point. Keeping all other properties fixed, an increase in $m$ also decreases the inducer concentration that is required to suppress repressor activity.
@@
While this role of induction has been widely discussed and used in experimental literature, theoretical studies leveraging this perspective for parametrization remain ongoing, with recent work demonstrating the utility of induction in modeling the dynamics of simple gene regulatory motifs [63, 57]. In particular, the scope of this approach has not been expanded to the more intricate regulatory networks standard in nature. We now incorporate this definition for the probability of active repressor into our model for gene expression dynamics in the three-gene toggle switch.

### v2

@@ Abstract
Control of transcription presides over a vast array of biological processes, including those mediated by gene regulatory circuits that exhibit multistability. Within these circuits, two- and three-gene network motifs are particularly critical to the repertoire of metabolic and developmental pathways. Theoretical models of these circuits, however, often vary parameters such as dissociation constants, transcription rates, and degradation rates without specifying precisely how these parameters are controlled biologically. In this study, we examine the role of effector molecules, which can alter the concentrations of the active transcription factors that control regulation, and are ubiquitous in regulatory processes across many biological settings. We specifically consider allosteric regulation in the context of extending the standard bistable switch to three-gene networks, and explore the rich multistable dynamics exhibited in these architectures as a function of effector concentrations. We then analyze how the dynamics evolve under various interpretations of regulatory circuit mechanics, underlying inducer activity, and perturbations thereof. Notably, the biological mechanism by which we model effector control over dual-function proteins transforms not only the phenotypic trend of dynamic tuning but also the set of available dynamic regimes. In this way, we determine key parameters and regulatory features that drive phenotypic decisions, and offer an experimentally tunable structure for encoding inducible multistable behavior arising from both single and dual-function allosteric transcription factors.
@@ I Introduction
Biological processes rely on intricate networks of gene interactions that together encode a multitude of possible cellular functions. Often, these genes produce proteins, called transcription factors, that alter the expression of other genes. Whether considering resource consumption in E. coli [59] or stages of animal development such as digit formation [75, 66, 92], vulval development [78, 37], or stem cell differentiation [81, 30, 39], such biological systems can precisely tune relevant gene interactions toward many possible distinct phenotypes. This inherent multistability is crucial to cell function, allowing signaling from the environment, metabolic resource constraints, or other sources to induce systems toward what at times can be dramatically different end states.
@@
Determining from experimental data alone the exact regulatory interactions that collectively drive multistable dynamics, however, is another matter. Indeed, the sensitivities of dynamics and steady state outcomes to both internal and external tuning are challenging to disentangle. To better understand these dynamics, we seek to explore motifs that establish the building blocks for multistability in higher-order networks.
@@
One regulatory motif that has emerged as an integral component of natural and synthetic genetic circuit design is the bistable toggle switch, in which each gene produces a protein that represses expression of the other. Study of the lysis vs. lysogeny decision in bacteriophage lambda over the 1960-80s established the significance of this motif experimentally [41, 62, 63, 27], and subsequent research has modeled such motifs extensively using Hill models [23, 10, 38]. As a two-state model, however, the bistable switch sheds only partial light on multistability in phenotypic expression.
Insights into the emergence of multistable dynamics in larger networks rely not only on the choice of model but also on the approach used. Existing literature on multistable cell fates during development, for example, has made significant strides by describing gene network dynamics within a Waddington landscape [85] and discussing the dynamic thresholds, or bifurcations, that can emerge as the underlying topology of an attractor system evolves [11, 12, 65, 69, 64]. These studies largely use geometric approaches that either coarse-grain representations to a set of network-free flow equations, or focus on systems that obey a gradient-like description. Given that many nontrivial networks contain feedback and cyclic interactions that break the monotonicity required for gradient-like behavior, however, other studies focus on the rotational effects that impact a system’s trajectory through the potential landscape [86, 6, 88]. These perspectives have been instrumental in characterizing a range of regulatory networks, often using abstract topological parametrizations and catastrophe theory. In this study, we build upon these rich traditions by directly linking bifurcation analysis of non-gradient multistable networks to experimentally accessible tuning knobs. In particular, we incorporate a concrete thermodynamic interpretation of induction to examine its explicit role in shaping multistable dynamics.
@@
In fact, the activity of transcription factors often depends on the presence or absence of allosteric effector molecules [25, 55, 56, 58, 14, 52, 53, 9, 26, 61]. When targeting a repressor, for example, an effector can act as an inducer by binding and stabilizing a protein in its inactive configuration, which suppresses its function and induces expression. Certain transcription factors are also known to have context-dependent behavior, acting as repressors at one promoter site along the DNA and as activators at another [2, 24]. Exploring different interpretations of the mechanistic role of effectors may then reveal critical differences in tuned expression that allow for complementary experimental investigations.
@@
Accordingly, our aim is to examine the role of the effector concentration as a system parameter, and by tuning this parameter alone, to directly observe corresponding dynamical shifts that occur. This framework offers a perspective with parametrizations that are accessible to the experimentalist while also inherent to cell activity itself. For example, researchers can directly input external non-metabolizable effectors to study the dynamic shifts observed in theory [50]. Cells also naturally leverage one or more effectors to tune dynamics, with two inducers driving cell fate determination in the lac operon and in cell differentiation during development, among other settings [70, 38]. It is important to recognize, of course, that tuning extracellular effector concentrations does not always exactly correspond to intracellular levels due to internal processes that can increase hypersensitivity, and that the internal concentrations of multiple distinct effectors can at times be codependent [43]. Nonetheless, effector concentration remains a valuable measure of a cell’s ability to regulate its gene expression dynamics, especially in light of the extensive experimental evidence emphasizing their significance as natural tuning knobs.
@@
To analyze the roles of effector molecules in the emergence of multistability, we specifically study extensions of the toggle switch to three genes. These motifs remain simple enough to visualize and physically interpret, yet also provide an additional layer of tunable gene interaction parameters that allows for more complex multistable dynamics – in this case, tristability and beyond. In exploring how a set of more than two genes couple to collectively decide among multiple possible cell fates, we also gain insights into how these interactions drive dynamics in larger and more complex networks.
@@
The three-gene toggle switch, shown in Figure 1, provides a powerful basis for examining the rise of tristability and higher-order multistability in a range of biological contexts. The motif is most directly useful in a synthetic setting, where dynamic shifts can be tuned precisely through experiments. However, the model can also help uncover how various sets of three (or more) key genes interact to generate the multistable cell fates observed in nature. For example, experimental work has identified sets of three “master regulator” genes that interact to drive protein expression toward three or more distinct cell fate phenotypes, such as in T helper cell differentiation [19] and stem cell differentiation [79].
@@
Recent work has also begun to build a comprehensive picture of stable expression in such three-gene networks [11, 16, 15]. Our study focuses instead on alternative approaches to induction in a Hill model for the three-gene toggle. Within this reframed context for tuning dynamics, we then investigate not only how the mechanisms of induction shift transcription factor activities and stable state expression levels, but also how the dynamic topologies of the resulting fixed point landscapes evolve.
@@
Figure 2: Expression of repressor 1 in the three-gene toggle switch, as described by a Hill function model. Equivalent definitions apply to expression of repressors 2 and 3. (A) Thermodynamic states, weights, and rates for expression of repressor 1 ($R_{1}$). The regulating repressors ($R_{2}$ and $R_{3}$) bind non-exclusively to the target promoter region to suppress gene transcription. Each repressor $R_{i}$ binds at the promoter region for repressor $R_{1}$ with affinity $K_{1i}$ and cooperativity $n$. (B) Thermodynamic states, weights, and rates for expression of repressor $R_{1}$ in the presence of inducers. Expression now depends on the active concentration of each repressor $R_{i}$, which is determined by a distinct inducer at concentration $c_{i}$, defining the probability of activity $p_{\text{act}}(c_{i})$.
We begin in Section II by defining our model for the three-gene toggle switch, and showing how it mathematically incorporates the regulatory activity of transcription factors as allosterically-inducible proteins. From this initial framework, Section III demonstrates the different dynamics that become possible as inducer concentrations rise. In Section IV we highlight how the thresholds, i.e., bifurcations, that separate different dynamic regimes shift when perturbed away from an initial assumed symmetry in gene interactions. Section V explores the impact of self-activation on the complexity of dynamics and the number of stable steady states possible for our three-gene switch. In particular, we find that the biological interpretation of dual repression and self-activation, whether arising from binding exclusivity or the effector’s role, can significantly affect the types of dynamics that unfold. In so doing, our work sheds light on the complexity of dynamics possible from simple modeling choices, and on just how precisely inducer concentrations must be tuned under different conditions for a system to access various dynamic regimes. The diverse range of input-output responses demonstrated across the different three-gene networks presented here motivates further study of how our results generalize to higher-order gene regulatory architectures.
@@ II The three-gene toggle switch
We first consider the most direct extension of the two-gene bistable toggle switch to three genes, as depicted in Fig. 1. Transcribing a given gene $g_{i}$ generates a protein $R_{i}$ called a transcription factor, which can in turn bind to the promoter region of either remaining gene. In this particular form of the model, the transcription factor prevents RNA polymerase (RNAP) from binding and initiating transcription, thereby repressing expression. We begin by discussing how to model the dynamics of this system of repressors before demonstrating the role of induction in governing how dynamical behavior evolves as a function of inducer concentration(s).
@@ II.1 The baseline model
This general form assumes that the dynamics of gene transcription, i.e., the production of mRNA transcripts, evolves at a timescale such that production can be measured equivalently from the concentration of output repressor $R_{i}$. The first term of Eqn. 1 characterizes protein production at a maximal rate $a_{i}$, and depends on the probability $p_{i}^{\text{expr}}$ that expression can occur. The second term accounts for protein degradation at a rate $1/\tau_{i}$.
Fig. 2(A) lists the possible regulatory states and corresponding statistical weights that contribute to the probability of expression $p_{i}^{\text{expr}}$. This approach to describing biological regulation draws from well-established statistical mechanical principles [35, 36, 1, 74, 5, 4, 7, 43, 84, 83, 82]. From this table, and equivalent definitions for expression of repressors 2 and 3, we define the dynamics of the corresponding three-gene circuit by the set of differential equations
@@
Each repressor $R_{j}$ binds non-exclusively with affinity $K_{ij}$ to the promoter region for the gene that expresses repressor $R_{i}$. This repression coefficient $K_{ij}$ reflects the characteristic concentration of repressor $R_{j}$ required to strongly regulate $R_{i}$’s expression. In other words, when $R_{j}=K_{ij}$, the produced concentration of repressor $R_{i}$ is half its maximum possible value.
Note that these equations describe transcription factor binding phenomenologically through a Hill function model with coefficient $n$ [33, 34]. This approach coarse-grains out the molecular detail of precise binding site occupancies found in a strictly thermodynamic model, instead representing its high cooperativity limit. The Hill coefficient $n$ thus does not necessarily correspond to the number of bound repressor molecules, but rather measures the sensitivity of output response to binding.
While there are advantages to the specificity of thermodynamic models [89], using them to evaluate stability profiles quickly becomes intractable in higher-dimensional networks, including those analyzed here. We therefore take a Hill model approach for the remainder of the discussion. Even without full knowledge of precise binding site occupancies, the Hill function offers a strong fit to empirical data for gene regulatory dynamics [23, 42]. The model also reflects the Hopfield barrier for the sharpness of input-output response [51], meaning that for a given coefficient $n$, the Hill model reflects the strongest, most sigmoidal input-output response possible for a system without energy expenditure. The approach thus allows us to investigate dynamics under the most input-sensitive model conditions, and to observe its maximum possible dynamic complexity.
We now have a model for the toggle switch that depends on various theoretically accessible parameters, specifically dissociation constants $K_{ij}$, production rates $a_{i}$, and degradation rates $1/\tau_{i}$. As pointed out previously, however, these parameters as written are not easily accessed experimentally, and are disconnected from the underlying biology that controls them within living cells. While the current model indicates that regulation is controlled by the total concentrations of transcription factors, these proteins are often themselves controlled by the allosteric binding of effector molecules. This implies that cells respond to changes in the number of active transcription factors.
@@ II.2 A statistical mechanical model for allosteric induction
Shifting to an allosteric description of gene regulation, we now argue formally that gene expression depends not on the total concentration of the regulating transcription factor, $\text{TF}_{\text{tot}}]$, but rather on the active concentration thereof, i.e.,
@@
where $p_{\text{act}}(c)$ is the probability that the transcription factor is active as a function of effector concentration $c$. This probability can be defined by the Monod-Wyman-Changeux (MWC) model [25, 55, 56, 58, 53, 9, 26, 61], which states that when an effector at concentration $c$ can bind allosterically at $m$ sites on a transcription factor, the probability of that transcription factor being active is
@@
Fig. 3(A) plots one possible probability curve $p_{\text{act}}(c)$ in the induction setting with $K_{I}<K_{A}$ for a fixed set of parameters. This function follows standard sigmoidal behavior, with a sufficient increase in inducer concentration rendering repressor inactive. Depending on how we define parameters, we can alter several key properties of the input-output response from this probability function. The leakiness represents the amount of activity at saturating concentrations of inducer, i.e., $p_{\text{act}}(c\rightarrow\infty)$, and the saturation conversely denotes the amount of activity in the absence of inducer ($p_{\text{act}}(c=0)$). The inducer concentration at which the probability is half maximal is denoted as the $EC_{50}$ value, and the sharpness of the input-output response curve at this inflection point rises with an increasing Hill coefficient $n$. The MWC parameters chosen therefore determine how effectively inducers suppress the activity of repressors and the sensitivity of the output response to an increase in inducer concentration.
@@
Figure 3: Repressor activity as a function of inducer concentration $c$, with $K_{I}<K_{A}$. (A) The probability of active repressor as a function of inducer concentration, with $m=2$, $\Delta\epsilon=4\>k_{B}T$, $K_{A}=150$ $\mu M$, and $K_{I}=5$ $\mu M$. The saturation and leakiness limits are denoted in blue and orange, respectively, and the $EC_{50}$, i.e., inducer concentration at which the probability is half maximal, is marked in purple. (B) Evolution of the probability for different values of $K_{I}$, with the curve in panel (A) shown again in green. The shift in probability spans the range of allowable $K_{I}$ values for the specified parameters up to the boundary at $K_{I}=6.77\>\mu M$, derived from Appendix A.
@@
Fig. 3(B) highlights several viable probability curves, defined by MWC parameters bounded to biologically permissible ranges derived from Eqns. 50 and 51 in Appendix A. We assume throughout that there are $m=2$ binding sites for inducers on repressors. For fixed $\beta\Delta\epsilon=4$ and $K_{A}=150$ $\mu M$, Fig. 3(B) then maps the probability of activity as a function of molar inducer concentration for increasing dissociation constant $K_{I}<K_{A}$. As the ratio $K_{A}/K_{I}$ decreases, approaching its lower bound (denoted in cyan as $K_{I}=6.77$ $\mu M$ in Fig. 3), the curve’s inflection point shifts toward higher inducer concentrations. This occurs because as $K_{I}$ rises and the inducer becomes less tightly bound to the inactive compared to the active repressor, a higher inducer concentration is necessary to sequester inactive repressors equivalently. Given that the choice of parameters in Fig. 3(A) is also within the biologically reasonable regime, we focus on these parameters for the remainder of our study, unless otherwise stated.
@@
Note that increasing the number of inducer binding sites $m$ for the target repressor protein would increase the cooperativity of inducer-repressor interactions, and thus increase the sensitivity of response to the presence of inducer. This would make the slope of the probability function steeper at the inflection point. Keeping all other properties fixed, an increase in $m$ also decreases the inducer concentration required to suppress repressor activity.
@@
While the role of induction has been widely discussed and used in experimental literature, theoretical studies leveraging this perspective for parametrization remain ongoing, with recent work demonstrating the utility of induction in modeling the dynamics of simple gene regulatory motifs [67, 89]. This approach, however, has not yet been expanded to the more intricate regulatory networks that are standard in nature. Toward this end, we now incorporate the MWC description for the probability of active repressor into our model for gene expression dynamics in the three-gene toggle switch.